OpenStudy (anonymous):

If a number is divided by eleven, the solution is two-fifths. What is the number?

OpenStudy (texaschic101):

x/11 = 2/5 -- multiply the equation by 11 x = 2/5 * 11 x = 22/5 check.. (22/5)/11 = 2/5 22/5 * 1/11 = 2/5 22/55 = 2/5 2/5 = 2/5 (correct) that number is 22/5

OpenStudy (anonymous):

4.4 or 22/5